OpenStudy (anonymous):

Please please please help What is the amplitude, period, and phase shift of f(x) = −3 cos(4x + π) + 6?

OpenStudy (anonymous):

you want to use the form \[f(x)=Acos(kx-\omega t+\phi _{0} )\] A:amplitude k:wave number x_position w:angular frequency t:time phi_0:phase

OpenStudy (anonymous):

do you mind walking through it or explaining it, because i don't really understand any of that... learning math online is very difficult on your own. @VeritasVosLiberabit

OpenStudy (anonymous):

sure. for one the easiest one to spot is the amplitude: -3 but since amplitude is always positive the amplitude is 3

OpenStudy (anonymous):

for the period I think you have to find x when 4x=2pi so the period is \[x=\frac{ 2\pi }{ 4 }\] and the phase is the tacked onto 4x which is pi in this case. Do you have an answer bank, because I'm not 100% on this and want to make sure its correct

OpenStudy (anonymous):

amplitude = 3; period = pi over two; phase shift: x = negative pi over four amplitude = −3; period = pi over two; phase shift: x = pi over four amplitude = −3; period = negative pi over two; phase shift: x = negative pi over four amplitude = 3; period = 2π; phase shift: x = negative pi over four

OpenStudy (anonymous):

the final answer should be \[A=3,T=\frac{ \pi }{ 2 },\phi _{0}=\pi \]

OpenStudy (anonymous):

ah ok so there is something off but close let me look

OpenStudy (anonymous):

looks like th closest one to your answer is A?

OpenStudy (anonymous):

I'm not sure where the -(pi)/4 is coming from though I need to think for a moment

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

hm maybe its not 4x=2pi for period but I think the answer is the same \[\frac{ 2\pi }{ T }=4\] is the period because that is what w is solve for T which is \[\frac{ \pi }{ 2 }\]

OpenStudy (anonymous):

so A must be right but the -(pi)/4 I'm not so sure about

OpenStudy (anonymous):

okay thank you for the help i appreciate it!! Do you mind helping with one more?

OpenStudy (anonymous):

sure make a new thread and link me

OpenStudy (anonymous):

Oh I think i see for the phase now you do \[4x+\pi=0\] and solve for x it seems